OpenStudy (anonymous):

Find the cube roots of 125(cos 288° + i sin 288°).

OpenStudy (anonymous):

Euler's formula?

OpenStudy (anonymous):

\[\cos{\theta}+i\sin{\theta} = e^{i\theta}\]

Parth (parthkohli):

No, a little more to that: De Moivre.

OpenStudy (anonymous):

which can be derived from Euler's, I was trying to get him to see that ;)

OpenStudy (anonymous):

There's a handy theorem called De Moivre's Theorem that will help with problems like this. It states that for a complex number \[\alpha(\cos \theta + i \sin \theta)^n = \alpha^n(\cos n\theta + i \sin n\theta)\] So you have to go backwards. Clearly the magnitude (a) will be 5, but what about the angle? You could just divide by 3, and get 96, but that's not the only option. You could also add 2pi and divide by 3, for 216, or subtract and divide, for -24.

Parth (parthkohli):

\[\sqrt[3]{125\cdot(\cos(288) + i\sin(288)} = \sqrt[3]{125} \cdot \sqrt[3]{\cos(288) + i\sin(288)} = 5\cdot \left(\cos(288) + i\sin(288)\right)^{1/3}\]

OpenStudy (anonymous):

How can I put that into my calculator with the "i" and pellet?

OpenStudy (anonymous):

Thanks for all the responses by the way

Parth (parthkohli):

You don't even need a calculator for this. lol

OpenStudy (anonymous):

i've believe i is 2nd zero on a 83 and cube root is in the math menu

Parth (parthkohli):

You just have to divide 288 by 3 somewhere. Is that so hard?

OpenStudy (anonymous):

2nd period, my bad, but yeah, you shouldn't really need a calculator

OpenStudy (anonymous):

(you should also try figuring out De Moivre's theorem from Euler's formula)

OpenStudy (anonymous):

I have no idea what any of this means. I have that formula but I dont use it? What?

Parth (parthkohli):

Just use this identity.\[(\cos\theta + i\sin\theta)^x = \cos(x\theta) + i\sin(x\theta)\]

Parth (parthkohli):

You have\[5\cdot (\cos (288^{\circ}) + i\sin (288^{\circ}))^{1/3}\]

OpenStudy (anonymous):

Why is it 5* and not 125?

Parth (parthkohli):

Scroll up...

OpenStudy (anonymous):

So I do 5* cos((1/2)288) + isin(1/2)288)? And I get -3.4573?

OpenStudy (anonymous):

Did I get it right @ParthKohli

Parth (parthkohli):

1/3*

Parth (parthkohli):

You shouldn't really use a calculator for this bro...

OpenStudy (anonymous):

A calculator might be necessary for computing the decimal component form, and convenient for getting the angles.

OpenStudy (anonymous):

I don't get how I get the root(s) of the original equation with that formula. Wouldn't the formula always give me a single number?

OpenStudy (anonymous):

Are my cube roots 96, 216, and -24?